What would be the number that, when added by 2, this sum multiplied by 3 and the total divided by 6 would result in 13?
Answer by addingup(3677)

Let me see if I understand:
First you add 2 to the number
x+2
the sum is multiplied times 3
(x+2)3 and the total divided by 6 would result in 13
[(x+2)3]/6 = 13
(3x+6)/6 = 13
3x+6 = 78
3x = 72
x = 72/3 = 24 is the number
------------------------------
check
(24+2)3/6 = 13
78/6 = 13 correct
:
John